OpenStudy (anonymous):

a married couple had a combined annual income of $63,000. the wife made $7,000 more than her husband. what was each of their incomes?

OpenStudy (anonymous):

x+7000=63000 should be your initial equation

OpenStudy (anonymous):

once x is isolated you need to divide by two to get the husbands salary

OpenStudy (anonymous):

then add 7,000 to the husbands salary to get the wives salary

OpenStudy (anonymous):

so the husband makes $28,000 per year and the wife makes $35,000 per year

OpenStudy (anonymous):

This question is invalid. Women make less money, as they should.

OpenStudy (anonymous):

\[ w+h = $63,000\\ w = h + $7,000 \]

OpenStudy (anonymous):

@bibby69, this is post divorce, post alimony income. Numbers are valid.

OpenStudy (anonymous):

Sub second equation into first one and get: \[ (h+$7,000) + h = $63,000 \]You can use this to find the husband income, then you can use second equation to find wife income.